OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The cost to purchase and install carpet in a 15 ft. by 16 ft. rectangular room was more than $6000. What was the price per sq. ft. to buy and install the carpet?

OpenStudy (anonymous):

The area of the room is A=l*w. A=15*16=240 sq. ft. We divide $6000 by the number of square feet to get 6000/240=$25/foot. Since it cost more than $6000 to install the carpet, it must have cost more that $25/foot to install.

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

No problem